Two cars start at the same point and travel in opposite directions. The first car travels 10 miles per hour faster than the second car. In 3 hours, the cars are 270 miles apart. What is the rate of the second car?
Answer by ankor@dixie-net.com(22740)

Two cars start at the same point and travel in opposite directions.
The first car travels 10 miles per hour faster than the second car.
In 3 hours, the cars are 270 miles apart.
What is the rate of the second car?
:
let s = the speed of the 2nd car
then
(s+10) = the speed of the 1st car
:
Write a distance equation: dist = time * speed
3s + 3s(+10) = 270
3s + 3s + 30 = 270
6s = 270 - 30
s = 240/6
s = 40 mph the 2nd car
then, obviously
10 + 40 = 50 mpn the 1st car
:
:
Check, find the actual dist each car travels
3*40 = 120 mi
3*50 = 150 mi
-----------------
total: 270 mi